### v1

# Estimating Parameter Fields in Multi-Physics PDEs from Scarce Measurements

Abstract
Parameterized partial differential equations (PDEs) underpin the mathematical modeling of complex systems in diverse domains, including engineering, healthcare, and physics. A central challenge in using PDEs for real-world applications is to accurately infer the parameters, particularly when the parameters exhibit non-linear and spatiotemporal variations. Existing parameter estimation methods, such as sparse identification and physics-informed neural networks (PINNs), struggle in such cases, especially with nonlinear dynamics, multiphysics interactions, or limited observations of the system response. To address these challenges, we introduce Neptune, a general-purpose method capable of inferring parameter fields from sparse measurements of system responses. Neptune employs independent coordinate neural networks to continuously represent each parameter field in physical space or in state variables. Across various physical and biomedical problems, where direct parameter measurements are prohibitively expensive or unattainable, Neptune significantly outperforms existing methods, achieving robust parameter estimation from as few as 50 observations, reducing parameter estimation errors by two orders of magnitude and dynamic response prediction errors by a factor of ten compared to PINNs. Furthermore, Neptune exhibits superior extrapolation capabilities, enabling accurate predictions in regimes beyond training data where PINN fail. By facilitating reliable and data-efficient parameter inference, Neptune promises broad transformative impacts in engineering, healthcare, and beyond.
Figure 1: Overview of Neptune, a generalizable method for inferring unknown parameter fields from sparse observations of system responses. Physical phenomena are often governed by known PDEs or coupled PDE systems with unknown key parameter fields. Neptune estimates these complex fields using extremely sparse physical response data through a two-stage process to ensure robustness. First, scalar parameters are estimated under an assumed homogeneous field approximation. Second, neural networks (referred to as $\mathcal{N}$) model the local variance of field parameters, along with the previous scalar estimations. Both stages iteratively solve the PDE system numerically, minimizing the error between predicted and observed dynamics via adjoint-based backpropagation. The trained method yields a reliable physical system model capable of predicting phenomena under varying environmental conditions (e.g., source terms, boundary conditions), enabling Extreme Event Resilience, Accurate Extrapolation, and Proactive Strategies as demonstrated in predicting battery thermal runaway behavior, characterizing flow in porous media, and detecting cardiac fibrosis via electrophysiological analysis.
Accurate modeling of physical phenomena is crucial in a wide range of scientific and engineering applications. Differential equations serve as a foundational framework for describing the dynamics of such systems, capturing processes that evolve over time and space. However, the parameters that govern the dynamics of such systems are often unknown or deviate from known values due to factors such as wear and tear, aging, or environmental variations, resulting in significant changes in the system’s behavior. For example, the thermal properties of aged batteries can evolve, increasing the risk of thermal runaway and fires in electric vehicles (EVs). Similarly, in biomedical contexts, alterations in the properties of cardiac tissue may lead to arrhythmias or atrial fibrillation, while variations in the flow properties within porous media can result in groundwater contamination or aquifer depletion. These examples underscore the critical need for parameter estimation, as failing to account for evolving parameters can lead to inaccurate predictions and severe consequences. Beyond accuracy, parameter estimation enables transformative applications, such as simulating future behavior, predicting breaking points, estimating remaining useful life in engineering systems, and creating personalized models for tailored healthcare diagnostics and treatments through more accurate digital twins.
Parameter estimation in multiphysics applications presents several interconnected challenges. The problem is inherently ill-posed, and in many real-world scenarios, direct measurement of parameters is impractical due to cost, inaccessibility, or physical constraints, leaving system responses as the only observable source of information. In addition to these difficulties, available data is often sparse, noisy, or incomplete, limiting the reliability of inference. Moreover, many relevant parameters are spatially distributed and field-dependent rather than scalar, and they are often tightly coupled across scales, making them difficult to disentangle or model independently.
Despite progress in parameter estimation methods, significant limitations remain. Traditional approaches such as finite element updating [1, 2, 3], Bayesian neural networks [4], least squares estimation [5, 6], Kalman filtering [7, 8], Gaussian processes [9, 10, 11], and sparse identification [12, 13] are effective in simplified, single-physics problems. However, they struggle with spatially and temporally varying parameters in nonlinear multiphysics systems. Karhunen–Loève expansions [14, 15, 16], while capable of estimating field parameters, have limited applicability to temporal dynamics [14, 15, 17] and rely on prior assumptions about parameter distributions [16].
In response to these limitations, recent machine learning based approaches such as PINNs [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], Neural Operators [33, 34, 35, 36, 37, 38, 39, 40, 41], and other deep surrogates [42, 43, 44] have gained attention as efficient alternatives for modeling complex, field-dependent parametric systems. However, a growing body of evidence [45] suggests that these methods often rely on weak numerical baselines and suffer from reporting biases, leading to overoptimistic claims of performance and generalization.
PINNs embed governing equations into neural network loss functions, enabling simultaneous learning of system dynamics and parameter fields [46, 47, 48, 49, 20, 50]. While effective in simple cases, they struggle to recover complex parameters from sparse or noisy data, especially in nonlinear, coupled multiphysics settings [51, 17, 52, 53, 54]. PINN-SR [13], which incorporates sparse regression principles similar to SINDy [12] for equation discovery, faces challenges in generalizing to systems with field parameters.
Neural Operators, such as DeepONet [33] and Fourier Neural Operators (FNO) [34], offer an alternative by learning solution operators directly from data and have been increasingly applied to inverse problems [55, 56, 57]. However, most methods first learn the forward map, then perform inverse estimation. Recent approaches such as Neural Inverse Operators, iFNO, and SC-FNO [58, 59, 55, 57] aim to bypass this structure but still require large, high-fidelity datasets [60, 58] with thousands of samples [55, 56]. Moreover, they remain sensitive to data distribution shifts [61] and require extensive hyperparameter tuning in complex problems [40, 61, 60].
Besides PINNs and neural operators, Neural ordinary differential equations (ODEs) [62, 63, 64, 65, 66, 67] model dynamics in the form of differential equations via neural networks. Coupled with adjoint methods [68, 63], they enable memory‑efficient parameter estimation. Yet the majority of existing work targets low‑dimensional ODEs with scalar parameters and known equations, without extending to PDE‑governed systems or spatially/temporally varying fields [69, 70, 71, 72].
Application Physics Parameters Measurements Parameter Extrapolated
Error Response Error
Battery Thermal & chemical $k$ (conduct.), $C_{p}$ (heat cap.) $T$ (temperature) 0.81%, 1.23% 0.36%
Porous Flow Darcy & dispersion $K$ (hydraulic conduct.) $u$ (concentration) 8.83% 0.69%
Cardiac Electro Aliev-Panfilov $D$ (tissue conduct.) $V$ (voltage) 3.01% 0.46%
Cell Migration Diffusion-reaction $\gamma$ (diffusion), $\lambda$ (growth) $\rho$ (density) - 6.40%
Table 1: Parameter estimation performance for a range of problems. Parameter error is defined as normalized mean absolute error, and extrapolated response error uses peak percentage error. Metrics are derived from approximately 125 measurements. Battery conductivity errors are reported for the $x$-$y$ direction only (similar in $z$). Cell migration (52 measurements, no parameter references) excluded parameter error, and the response error is obtained from the training region.
Summary of the Neptune results in the context of accuracy for a range of models Problem Err. (N-0%) Err. (N-1%) Err. (N-10%) Description Battery 0.01 ($\pm$0.006) 0.05 ($\pm$0.009) 0.15 ($\pm$0.038) $k(T)$ Porous Media Flow 0.089 ($\pm$0.019) 0.10 ($\pm$0.016) 0.21 ($\pm$0.073) $K(x,y)$ Cardiac Electrophysiology 0.0030 ($\pm$0.00082) 0.016 ($\pm$0.0060) 0.024 ($\pm$0.0052) $D(x,y)$
Table 2: The error is defined as the average absolute error of the estimated parameter compared to the reference. The values in the parentheses denote the noise levels (e.g., noise-free 0%, 1%, and 10%). The samples used in the above cases are around 128. In the battery problem, only the thermal conductivity $k$ estimation is displayed for simplicity. Absolute error is used instead of percentage error, as values near zero can distort the error assessment.
To address these challenges, this article introduces Neptune (Neural Estimation of Parameters in mulTi-physics PDEs Under sparse obSErvations), a robust and generalizable method for field parameter estimation (see Figure 1). Neptune leverages governing equations and integrates numerical solvers with neural network-based parameter inference, achieving both high parameter estimation accuracy and data efficiency, even under extremely sparse and noisy observation settings. More importantly, it employs a two-stage estimation strategy, first recovering coarse scalar parameters, then refining local variations via neural networks, enabling robust and scalable estimation of complex, multiple field parameters in nonlinear multiphysics systems. In addition, Neptune demonstrates strong generalization and extrapolation capabilities, delivering accurate predictions beyond training regions and across diverse scientific domains, all without assuming parameter distributions or full system knowledge. To support this, governing PDEs are discretized into systems of ODEs, enabling efficient adjoint-based optimization through the numerical solver. Meanwhile, the unknown parameter fields are estimated without any prior assumptions about their distributions, further broadening the applicability of Neptune to multiphysics and coupled systems.
In the following sections, we demonstrate the versatility and efficacy of Neptune across various scientific domains, including cardiac electrophysiology, cell migration and proliferation, thermal-chemical battery systems, and flow dynamics in porous media. Through these applications, we highlight the method’s ability to deliver robust, accurate results under data-scarce and noisy conditions, its resilience to data noise, and its adaptability to complex scenarios using real-world datasets. These qualities underscore Neptune’s transformative potential in advancing parameter estimation and predictive modeling for real-world systems. Quantitative results in Tables 1 and 2 further underscore these strengths, illustrating Neptune’s superior performance in parameter estimation across diverse settings.
Figure 2: Parameter estimation in cardiac electrophysiology. a. The training region with Gaussian noise during measurement, and the estimated forward inference of $V$. The predictions are compared between Neptune and PINN. b. The parameter estimation error under varying training data sizes. A dashed line is used as a reference beyond 1,000 data points to indicate that high performance has already been achieved, with no further improvement expected. c. The parameter estimation distribution comparison between Neptune and PINN, for two representative cases of $D$. e. The estimated forward response $V$ comparisons for case 1. When performing temporal extrapolation, PINN predicts values that do not adhere to physics laws.
I Cardiac Electrophysiology
Cardiac electrophysiology (EP)  [73, 74, 75, 51] emphasizes the crucial coupling between cardiac tissue properties and the generation as well as propagation of electrical signals within the heart.
Known Physics: The canine ventricular Aliev-Panfilov model is utilized in this section, with the coupled equations [51] defined in the following.
$\frac{\partial V}{\partial t}=\nabla(D\nabla V)-k_{0}V(V-a)(V-1)-VW$ (1)
$\frac{\partial W}{\partial t}=\big{(}\epsilon+\frac{\mu_{1}W}{V+\mu_{2}}\big{)}\big{(}-W-k_{0}V(V-b-1)\big{)}$ (2)
This model is commonly used to identify tissue heterogeneities, particularly from in silico data [76, 77, 51], with significant clinical potential for detecting fibrosis and other localized pathologies linked to arrhythmias and atrial fibrillation [51, 77]. The diffusion tensor $D$, the target parameter, determines propagation speed, with $V$ as the measurable voltage and $W$ as an intermediate state variable. The PDEs describe the evolution of electrical potential and ionic currents in cardiac tissue, incorporating known properties such as tissue conductivity and reaction-diffusion dynamics.
Unknown Parameters: The key parameter, tissue conductivity (proportional to the diffusion tensor $D$), varies spatially and is challenging to measure directly, requiring estimation for accurate modeling. For simplicity, the estimation focuses on $D$. Details of all known physical parameters and conditions relevant to this problem, along with those for other cases discussed in the article, can be found in Supplementary Note 1.
Problem Setup: A 2D slab of cardiac tissue is employed ($1cm\times 1cm$) as the spatial domain and aims to recover the heterogeneous diffusion parameter $D(x,y)$. Healthy tissue is represented by $D=0.1mm^{2}/TU$, while fibrosis tissue has $D=0.02mm^{2}/TU$ five times smaller. $1TU$ is approximately $13ms$ [51]. During training, Gaussian noise is added to the training data ($V$ measurements) to mimic real-world situations. The voltage response over time (from 0 to 40 $TU$) at a randomly selected spatial point, including noise, is shown in Fig. 2a.
Measurements (with details and comparison to PINN): Additionally, a PINN model [51] is compared to Neptune for joint parameter estimation and dynamic prediction. PINN requires at least 8,000 spatiotemporal samples on a finer $\times100$ mesh, while Neptune achieves comparable results with only 62 sparse observations on a $\times50$ mesh. Figure 2b compares Neptune and PINN in parameter estimation across different training data sizes. Neptune achieves highly accurate and stable parameter estimates with as few as 1,000 samples, while PINN consistently shows higher errors at all data levels. Furthermore, Neptune significantly outperforms PINN even with noisy data, demonstrating superior accuracy under challenging conditions and limited training samples.
Analysis: Figure 2c evaluates the parameter estimation performance of Neptune and PINN across two cases with different parameter distributions. The first column shows the reference distributions, used to generate ground truth data for training. In the second column, Neptune successfully recovers the parameters in both cases. By contrast, in the fourth column, PINN struggles significantly, even with 50 times more data (50,000 vs. 1,000 points), resulting in a mean absolute error (MAE) that is three orders of magnitude larger. In Case 1, PINN performs worse overall, and in Case 2, it fails to capture the random distribution of $D$.
Figure 2d compares state variable $V$ predictions for Case 1 between Neptune and PINN. While PINN performs reasonably well within the training region (first row), its peak error is an order of magnitude higher than Neptune’s. For temporal extrapolation (second row), PINN shows significantly larger errors due to its inaccurate estimation of $D$. As shown in Fig. 2a, Neptune consistently adheres to the underlying physics, providing robust predictions even beyond the training region. It achieves accurate temporal extrapolations beyond 40TUTU, with an MAE of 0.00035, which is over two orders of magnitude lower than PINN’s MAE of 0.066. Moreover, PINN frequently predicts negative $V$ values at multiple timesteps, further indicating its limitations. Additional results for parameter estimation and forward inference under extremely sparse data conditions are detailed in Supplementary Note 2.
Implications: The ability of Neptune to accurately estimate spatially heterogeneous parameters and predict state variables even under sparse, noisy conditions showcases its superiority over traditional methods like PINN. This advancement holds great promise for improving the diagnosis and modeling of cardiac pathologies, such as fibrosis and arrhythmias, where tissue heterogeneities play a critical role in understanding electrical signal propagation.
Figure 3: Field parameter estimation in thermal runaway problems. a. A prismatic battery, a cylinder battery, and random temperature measurements at the battery surface. Cross-sections AAA-A and BBB-B represent the cross-plane and the in-plane directions, respectively. b. The estimated three field parameters $C_{p},k(y-z)$, and $k(x)$ for the prismatic battery. c. Predicted thermal runaway behaviors with the estimated parameters. The initial temperature of the battery is 25°$C$ (or 298.15KK), and the boundary condition (ambient temperature) is 423.15KK. After accurately estimating the parameters, the model can correctly predict the thermal and chemical responses. d. Evaluation of temperature distribution for the prismatic battery at two typical times as marked in graph c. e. Predicted thermal runaway behaviors for the cylinder battery. f. Evaluation of temperature distribution for the cylinder battery.
II Thermal Runaway in Batteries
Modern lithium-ion batteries are designed to achieve high energy density, extended operational durations, and rapid charging capabilities, catering to the power requirements of contemporary electric vehicles (EVs). These high-performance batteries provide substantial benefits but also pose safety challenges, particularly the risk of thermal runaway (TR), which in extreme cases can result in fire accidents [78]. To address these risks, a thorough understanding of the underlying physics is crucial.
Known Physics: TR is often governed by well-established thermal reactions and chemical kinetics, which are mathematically described by the following equations.
$\rho C_{p}\frac{\partial T}{\partial t}=\nabla(k\nabla T)+\dot{Q}_{exo}$ (3)
$\dot{Q}_{exo}=HW\left(-\frac{\partial c}{\partial t}\right)$ (4)
$\frac{\partial c}{\partial t}=-A_{e}exp\left(-\frac{E}{R_{c}T}\right)c$ (5)
$T$ and $c$ represent the battery temperature and concentration of reacting materials. Other variables such as energy release $\dot{Q}_{exo}$, heat release $H$, active material content $W$, activation energy $E$, gas constant $R_{c}$, and porosity $\phi$ are described in more detail in Supplementary Note 1.
Unknown Parameters: The parameters of interest  [79, 80, 81] are temperature-dependent heat capacity $C_{p}$, thermal conductivity $k$, and a scalar parameter called the pre-exponential factor $A_{e}$. For prismatic batteries, thermal conductivity is divided into cross-plane ($k_{z}$) and in-plane ($k_{xy}$) components, whereas for cylindrical batteries, it is characterized by radial $k_{r}$ and longitudinal $k_{z}$ components. The pre-exponential factor $A_{e}$ is specifically introduced to illustrate the generalizability of the method.
These parameters are difficult to measure directly and are expected to behave as field variables influenced by the battery’s internal states, such as temperature and other dynamically changing properties during operation. Additionally, properties like thermal conductivity and heat capacity can exhibit significant variability due to mass production inconsistencies and further evolve over their service life from aging effects [82] and environmental influences [83]. This variability makes these parameters difficult to measure directly, but essential for understanding battery performance and predicting the likelihood of TR incidents.
Problem Setup: A standard prismatic battery ($100mm\times 180mm\times 32mm$) and a cylindrical battery ($66mm$ height, $28mm$ diameter) (see Fig. 3a) are modeled to simulate TR behavior [84]. The initial condition is room temperature, with constant heat convection as the boundary condition (BC), and high external heat applied to simulate overheating.
Measurements: Considering real-life constraints, only surface temperatures can be measured, and only at limited locations. For training, we take just three measurement locations, with 15 spatial-temporal measurements each, during the initial 100 minutes before the irreversible temperature rise at around 150 minutes in TR.
Analysis: Despite this limitation, Neptune accurately estimates the field parameters ($C_{p},k_{x}$, and $k_{z}$). For the prismatic battery, the estimated parameters (Fig. 3b) show MAE of 1.88, 0.052, and 0.001, compared to reference values of 1400, 13, and 0.8, respectively. For the cylindrical battery, the scalar parameter $A_{e}$ is jointly estimated, with an estimated value of $5.19\times 10^{25}$ showing a low percentage error of 0.97% compared to the reference $A_{e}=5.4\times 10^{25}$. Similar accuracy is achieved for other parameters, with results detailed in Supplementary Note 2. Furthermore, Neptune also demonstrates robustness under noisy measurements; for the prismatic battery, results for noise levels of 1% and 10% are detailed in Table 1.
Implications: Importantly, accurate parameter estimation enables the prediction of temperature changes of potentially ongoing TR. By integrating the parameters into the PDE system, predictions are extended over 0-270 minutes. Figure 3c shows the temporal predictions of average temperature and concentration ($c$) profiles, closely matching the ground truth. Besides, inner temperature distributions can also be evaluated. In Fig. 3d, representative results are displayed for cross-sections AAA-A and BBB-B (Fig. 3a) at two key stages, before and during TR, where temperature profiles are accurately predicted with minimal errors.
Additionally, accurate parameter estimation using Neptune enables model customization to simulate battery behavior under diverse initial and boundary conditions (illustrated in Supplementary Note 2). This capability supports predictive modeling for various scenarios, including environmental changes, manufacturing inconsistencies, and operational extremes, which are crucial for enhancing battery safety and performance. Unlike existing methods such as PINN, which often fail due to reliance on specific initial and boundary conditions, Neptune is independent of these constraints, ensuring consistent and reliable performance across scenarios. This robustness unlocks new opportunities for designing safer, more efficient batteries and evaluating their behavior over extended lifetimes or under extreme conditions.
Figure 4: Field parameter estimation for flow in porous media. a. A demonstration of flow phenomena, and random measurements are taken as the ground truth for model training. b. The estimated hydraulic conductivity $K$ compared to the reference $K$. c. The estimated field parameter $K$ and its estimation error compared to the reference, for different numbers of training samples. The magnitudes are normalized. d. The particle concentration predictions (state responses) at different times. The predictions align well with the ground truth, even for extrapolation at $t=20$ minutes. e. The parameter estimation performance compared to different numbers of training samples. Error bands are derived from standard deviations from multiple estimations and represent the variability in the performance. f. The response prediction performance compared to different numbers of training samples.
III Flow in Porous Media
The flow of solutes in porous media is a fundamental phenomenon observed in various natural systems [85, 86, 87, 88, 89, 90, 91]. This includes critical environmental concerns such as contaminant transport in water and the greenhouse effect caused by carbon emissions.
Known Physics: As an illustrative example, we focus on subsurface transport, a problem frequently studied in recent literature [92, 17, 53]. This process is governed by coupled PDEs: the time-dependent advection-dispersion equation (ADE) and Darcy’s law. These equations describe time-dependent transport behaviors and the physics of fluid flow in porous media, which are key to understanding and predicting such systems.
$\frac{\partial u}{\partial t}+\nabla\cdot[\mathbf{v}u]=\nabla\cdot[\mathbf{D}\nabla u]$ (6)
$\nabla\cdot[K\nabla h]=0$ (7)
$\mathbf{v}=-K\nabla h/\phi$ (8)
where state variable $u$ represents the particle concentration field. Variables such as average pore velocity $\mathbf{v}$, dispersion coefficient $\mathbf{D}$, hydraulic head $h$, and porosity $\phi$ are described in more detail in Supplementary Note 1.
Unknown Parameters: In this system, solute movement through porous media varies due to environmental conditions, human activities, and geological processes, making accurate parameter estimation of hydraulic conductivity $K(x,y)$ crucial for understanding these phenomena and addressing these pressing issues  [93, 17, 53, 94]. However, $K$ is often difficult to measure directly, whereas particle concentration fields, depending on the system and measurement tools, can be relatively easier to retrieve.
Problem Setup: A 2D computational domain of $L_{x}=1m$ and $L_{y}=0.5m$ is defined [92], as shown in Fig.4a. Using just 32 random spatial-temporal concentration measurements $u$ over 0–16 minutes, the field parameter $K$ can be robustly estimated. Since the right-hand side of equation 7 is 0, multiple solutions exist, so the estimated $K$ is normalized and compared to the normalized reference (see Fig. 4b).
Measurements: Figure 4c shows the normalized $K$ estimation for different sample sizes ($N=$ 32, 64, 256). With fewer samples, the data may not fully capture the response, leading to larger prediction errors. However, even with just 32 points, Neptune already captures the key features of the parameter’s distribution. Increasing the number of samples enhances the resolution of predictions and further reduces estimation errors.
Analysis: Figure 4d illustrates the evolution of $u$ predictions over time. The first and second rows show contour predictions during the training region (at $t=5$ minutes and $t=10$ minutes) with low error and strong alignment with the ground truth. Notably, temporal extrapolation up to 20 minutes, beyond the 16-minute training region, accurately captures the concentration $u$, as shown in the third row of Fig. 4d.
To demonstrate the robustness of Neptune, parameter estimation is repeated with varying sample sizes. Figure 4e shows the parameter estimation performance, while Fig. 4f displays the corresponding response prediction accuracy, with errors significantly reduced at 64 samples. Ablation studies on neural network sizes, detailed in Supplementary Note 4, further validate the model’s robustness and generalizability. Additionally, a spline-based parameter fitting method was tested as a non-neural network strategy but showed lower performance than Neptune (around 2 times larger mean errors in both the parameter estimation and response prediction). The spline-based parameter fitting setups and results can be found in Supplementary Note 6.
Implications: Accurately estimating the hydraulic conductivity field $K$ and predicting solute transport dynamics with Neptune enables robust modeling of subsurface transport, even with sparse and noisy measurements. This capability is critical for addressing environmental challenges such as contaminant transport and groundwater management, where accurate parameter estimation and temporal predictions are essential for effective decision-making.
Figure 5: Parameter estimation in cell migration and proliferation. a. Parameter estimation results and comparisons. Neptune assumes all three parameters are density-dependent. The results show an increasing $\gamma$ value as density increases, while the other two parameter estimations exhibit minimal variation and can be considered as constants. This aligns well with assumptions from previous research [95, 96]. b. Density response predictions. Using the estimated field parameters, the density response at four different timesteps is displayed. Compared to results using the reference method (PINN-SR) and the scalar parameters setting, Neptune more accurately matches the experimental data at 36-hour and 48-hour, particularly in the regions with higher density values on both sides.
IV Cell Migration and Proliferation
Application and Known Physics: This section presents an experimental study of the diffusion (cell migration) and reaction (cell proliferation) process, which is governed by a known form of PDE [97, 98, 13].
$\frac{\partial\rho}{\partial t}=\gamma\frac{\partial^{2}\rho}{\partial x^{2}}+\lambda_{1}\rho-\lambda_{2}\rho^{2}$ (9)
where parameters such as cell diffusivity $\gamma$ and proliferation rates $\lambda_{1}$ and $\lambda_{2}$ define the characteristics of cell migration and proliferation.
Unknown Parameters: These parameters are unknown and require estimation from cell density $\rho$ observations. According to the Porous-Fisher model [95, 96], cell diffusivity $\gamma$ varies with density (Fig. 5b), increasing with higher cell density, while proliferation rates $\lambda_{1}$ and $\lambda_{2}$ are treated as constants. However, to demonstrate Neptune’s generalizability, all 3 parameters are modeled as density-dependent fields.
Problem Setup: The experimental data were originally extracted from high-resolution imaging collected in previous research on in vitro cell migration (scratch) assays [99], and were further analyzed [13] aimed at discovering the governing equation. The preprocessed data describes the cell density distributions at 38 evenly distributed spatial points in one dimension and only at 5 time steps ($0h,12h,24h,36h,48h$). The spatial domain ranges from 0 to 1900 $\mu m$. A detailed description of the experiment setup and data preprocessing can be found in the previous research [99, 13]. This spatial-temporal experimental data is extremely sparse and noisy.
Measurements: Given the initial condition at 38 points from the experimental data ($t=0h$) and Neumann boundary conditions $\partial\rho(x=0)/\partial x=0$ and $\partial\rho(x=1900)/\partial x=0$, we train Neptune with only 152 cell density data points over 38 spatial points and 4 timesteps. In the same setup, we compared our Neptune with PINN-SR [13] (i.e., PINN with sparse regression), for parameter estimation and response prediction performance.
Analysis: As shown in the first graph of Fig. 5a, Neptune estimates an increasing $\gamma$ with cell density, consistent with previous research [97, 98]. Field estimations also suggest $\lambda_{1}$ and $\lambda_{2}$ are likely scalars. We also assume space-dependent parameters for estimation: results showing parameter distributions with high (peak) values at two boundaries, demonstrating that the field parameters are density (or state) dependent. In the third graph, while the field estimation for $\lambda_{2}$ shows large values at low densities and stabilizes as density increases, this behavior is primarily due to highly noisy experimental data in regions with low cell numbers, rather than being attributed to the nature of the field parameter itself. In contrast, PINN-SR, limited by its methodology, cannot estimate field parameters, and its scalar estimations show slightly different combinations but remain within a reasonable range. This limitation is evident at $12h$ (first graph) and $48h$ (last graph), where it fails to capture enhanced cell spreading.
Accurate field parameter estimation excels in predicting cell density distributions over time, as shown in Fig. 5b. For such sparse and noisy data, PINN-SR achieves a median error of $6.53\times 10^{-5}$. In contrast, Neptune, assuming density-dependent parameters, achieves the highest accuracy with an error of $4.71\times 10^{-5}$. Across experimental data with varying initial cell densities [13], Neptune achieves 28-47% lower median errors. Additional results are detailed in Supplementary Note 2.
Implications: The findings demonstrate that Neptune not only provides more accurate parameter estimations under sparse and noisy conditions but also captures critical density-dependent dynamics that other methods, like PINN-SR, cannot. This capability enables improved predictions of biological phenomena such as cell migration and proliferation, offering new avenues for understanding and modeling reaction-diffusion systems in complex and noisy experimental scenarios.
V Discussion
In this paper, we presented Neptune for accurate field parameter estimation under scarce observation data, demonstrated by various PDE problems in emerging fields where physical parameters are unknown and important. Specifically, we estimated the thermal conductivity and heat capacity in coupled thermal problems, enabling accurate TR prediction. In porous media flow, spatially varying hydraulic conductivity was recovered, improving the modeling of phenomena such as contaminant transport. For cardiac EP, tissue electrical conductivity was inferred solely from voltage measurements, offering potential for detecting fibrosis and other localized pathologies. Finally, the cell migration and proliferation case demonstrates effective parameter estimation under real-world noisy and sparse data conditions. It also highlights the method’s capability in discovering governing equations and its generalizability in estimating scalar variables.
In terms of utility, Neptune demonstrates its generalizability and robustness to estimate multiple parameters simultaneously across various dependencies (spatial, state, and even time), each exhibits distinct patterns. The proposed two-stage learning strategy further enhances effectiveness by handling parameters across different scales. By first estimating scalar magnitudes, it significantly reduces the search space, allowing neural networks to focus on learning local variations within the parameter field. Furthermore, we extend this discussion to demonstrate the superior performance of our strategy compared to other training approaches, including: (1) estimating only scalars, (2) directly estimating each field parameter in one step using separate neural networks; (3) estimating all field parameters within a single neural network; and (4) jointly estimating scalar magnitudes and field variations in a single step. Detailed comparisons are provided in Supplementary Note 5. Compared to other fitting methods, Neptune significantly outperforms spline-based approaches in capturing complex parameter distributions. In addition, the use of FDM makes the method easily adaptable to a wide range of PDEs. While less effective for handling shock waves or large discontinuities, it offers favorable computational efficiency and supports gradient-based optimization.
Compared to sparse identification methods like PINN-SR, Neptune models full field parameter distributions rather than scalars, achieving 27–47% improvement in forward response accuracy. This methodological shift enables more detailed modeling of the underlying physical processes, leading to improved predictive accuracy. Additionally, compared to PINN, our framework offers stronger temporal extrapolation, better adherence to physical laws, and greater adaptability to boundary and initial condition changes. It also requires less data, making it ideal when measurements are limited or costly. However, inverse problems like parameter estimation remain ill-posed and prone to local minima, especially with complex fields or limited data. These challenges can be mitigated by using informed initial guesses and incorporating prior knowledge, such as parameter magnitudes, smoothness penalties, and strategic data selection. Future work may extend the method to more general finite element frameworks and explore its application to modeling and parameter estimation in chaotic systems.
VI Methods
Problem Setup and differentiable PDE solver We consider a general time-dependent PDE of the form encompassing a range of PDE-governed systems studied in this work:
$f\left(\frac{\partial\mathcal{S}}{\partial t},\mathcal{S},\nabla\mathcal{S},\nabla^{2}\mathcal{S},t;\mathcal{P},\ldots\right)=0,$ (10)
where $\mathcal{S}=\mathcal{S}(\mathcal{X},t)$ is the state variable defined over spatial domain $\mathcal{X}$ (e.g., $x$, $y$, or $z$) and time interval $t\in[0,T]$. The unknown parameters $\mathcal{P}=[p_{1},p_{2},\dots]$ represent latent or field-dependent quantities of interest. The terms $\nabla\mathcal{S}$ and $\nabla^{2}\mathcal{S}$ denote the spatial gradient and Laplacian of the state variable, respectively. Neptune directly solves PDE systems by applying FDM for spatial discretization and using differentiable ODE solvers for time integration. This approach ensures efficient and accurate forward simulations while enabling gradient computation through the solver via the adjoint method, significantly enhancing training speed. Using evenly spaced mesh nodes, the second derivative in the $x$-direction at a point $x=x_{i}$ is approximated by the central difference scheme [100]:
$\frac{\partial^{2}\mathcal{S}(\mathcal{X}_{i},t)}{\partial\mathcal{X}^{2}}=\frac{\mathcal{S}_{i-1}-2\mathcal{S}_{i}+\mathcal{S}_{i+1}}{\Delta\mathcal{X}^{2}}+\mathcal{O}(\Delta\mathcal{X}^{2})$ (11)
where $\Delta\mathcal{X}$ is the mesh spacing and $\mathcal{S}_{i}$ is the discrete state value at $\mathcal{X}_{i}$. The term $\mathcal{O}(\Delta\mathcal{X}^{2})$ denotes the truncation error, indicating second-order spatial accuracy.
In most scenarios considered in this study, the spatial dimensions of the state variables are organized into a 1D vector form for computational efficiency. The state variable and discretization matrices are structured as $u\in\mathbb{R}^{n\times 1}$ and $A_{j}\in\mathbb{R}^{n\times n}$, where $n=n_{x}\times n_{y}\times n_{z}$ denotes the total number of spatial nodes in the discretized 3D domain. Spatial derivatives of various orders are explicitly formulated for each direction, as shown in equation 12:
$\frac{\partial^{j}\mathcal{S}}{\partial\mathcal{X}^{j}}=A_{j}\mathcal{S},$ (12)
where $j$ represents the derivative order and the matrix $A_{j}$ numerically approximates the corresponding spatial derivative.
Substituting the finite difference discretizations into the general PDE form yields the semi-discrete system:
$f\left(\frac{d\mathcal{S}}{dt},\mathcal{S},A_{1}\mathcal{S},A_{2}\mathcal{S},t;\mathcal{P},\ldots\right)=0,$ (13)
where $A_{1}$ and $A_{2}$ denote the numerical approximations of the first- and second-order spatial derivatives, respectively. Boundary conditions are carefully handled during discretization. Two types, Dirichlet and Neumann, are considered, expressed as $\mathcal{S}(\mathcal{X})=\beta$ or $\frac{\partial\mathcal{S}(\mathcal{X})}{\partial\mathcal{X}}=\beta$, where $\beta$ is the prescribed coefficient. Boundary conditions are implemented by modifying the discretization matrices in equation 12, following standard approaches [101, 102]. After discretization, the PDE system is reformulated as a set of ODEs and solved using Runge–Kutta methods [103], with dynamics extracted at observation times.
Field parameter modeling with neural networks Neptune models a unknown field parameter $p$ as:
$\mathcal{P}=\tilde{\mathcal{P}}\big{(}1+\mathcal{N}_{\bm{\theta}}(\mathcal{X}\text{ or }\mathcal{S})\big{)}$ (14)
$\mathcal{N}_{\bm{\theta}}=f_{m}\circ f_{m-1}\circ\cdots\circ f_{1}(\mathcal{X}\text{ or }\mathcal{S})$ (15)
where $\tilde{p}$ denotes a learnable scalar and $\mathcal{N}$ represents the neural network with $m$ layers and learnable weights as $\bm{\theta}$. A detailed architecture with skip connections can be referred to Fig. 6. $f$ denotes a neural layer and $\circ$ represents function composition. The input to $\mathcal{N}$ can be discretized values of space $\mathcal{X}$ or state $\mathcal{S}$, and are properly scaled (i.e., -1 to 1 from $\mathcal{X}$ or observed $\mathcal{S}$). Leaky ReLU activation functions are used in all layers except the final one, which omits activation functions. For multiple field parameters $\tilde{\mathcal{P}}=[p_{1},p_{2},\dots]$, Neptune simply employs different scalars $\tilde{p}_{i}$ and networks $\mathcal{N}_{{\bm{\theta}}_{i}}$ for parallel modeling, where $i=1,2,\dots$
Figure 6: The schematic diagram of the two-stage field parameter estimation strategy and the neural network architecture. a. The feed-forward neural network. Left: overall architecture. Right: residual block structure. For flow and cardiac electrophysiology problems, the input to the network is the spatial coordinate $\mathcal{S}$. For battery thermal runaway and cell migration, the input is the state variable $\mathcal{X}$ (e.g., $T$ or $\rho$), with fewer layers used. The network output $\mathcal{N}_{{\bm{\theta}}}$ is scaled to the physical parameter as $\tilde{\mathcal{P}}\big{(}1+\mathcal{N}_{{\bm{\theta}}}\big{)}$. The reference value $\tilde{\mathcal{P}}$ is application-specific and determined via a two-step training strategy. b, Parameter estimation diagram. The left part illustrates the two-stage estimation for scalar and field parameters, while the right part shows the numerical solving and optimization procedure.
Two-stage parameter estimation strategy Neptune employs a two-stage learning, first optimizing $\tilde{\mathcal{P}}$ and then training $\mathcal{N}_{\bm{\theta}}$. In the first stage (see the upper row of Fig. 6), random scalars $\tilde{\mathcal{P}}$ are initialized and incorporated into the discretized governing equations to numerically solve for the dynamic response $\hat{u}$. This is equivalent to setting $\bm{\theta}=\mathbf{0}$ in equation 14. Those learnable scalars are optimized using the adjoint method [62] (detailed later in the section) by minimizing the error between $\hat{\mathcal{S}}$ and the observed measurements $u$. The second stage fixes the converged scalars and only optimizes the neural network weights $\bm{\theta}$, aiming to estimate local variations in a reduced search domain (see the lower row of Fig. 6).
During the optimization, scalars $\tilde{\mathcal{P}}$ act as a physics-based prior and implicit regularizer, ensuring early physical stability of training and providing a well-scaled baseline for convenient expansion to field parameters in the second stage. Even if $\tilde{\mathcal{P}}$ converges with high error, its combination with the PDE remains physically meaningful and stable during optimization.
In the second stage, the neural network acts primarily as a correction term, analogous to residual learning architectures such as ResNet [104]. Initially, the fixed scalar $\tilde{\mathcal{P}}$ constrains the neural network to explore only perturbations around a physically meaningful baseline. This constraint stabilizes gradient flow, prevents nonphysical field magnitudes, preserves the well-posedness of the underlying PDE, and generally improves training stability. Meanwhile, in later stages of training, the neural network is allowed unbounded exploration (in most cases, no activation is applied to the final layer), enabling it to capture more complex behaviors beyond the initial physics-based prior.
Compared to directly learning $\tilde{\mathcal{P}}$ with a neural network, the first-stage scalar learning significantly accelerates training, as numerically solving the PDE with field parameter modeling is several times more expensive, depending on model complexity. It also reduces the search space and greatly stabilizes the second-stage learning of $\mathcal{N}_{\bm{\theta}}$.
In addition, by anchoring the field prediction around a physically meaningful baseline, this strategy improves generalization and robustness during training. The gradient flow with respect to $\bm{\theta}$ inherits better conditioning from the scalar optimization step, further enhancing convergence efficiency and training stability. Mathematically, the gradient with respect to the scalar parameter in the first stage is given by
$\frac{dL}{d\tilde{\mathcal{P}}}=\frac{dL}{d\mathcal{P}}\frac{d\mathcal{P}}{d\tilde{\mathcal{P}}}=\frac{dL}{d\mathcal{P}},$ (16)
where $d\mathcal{P}/d\tilde{\mathcal{P}}=1$ because $\mathcal{P}=\tilde{\mathcal{P}}$ in state one.
In the second stage, following the reparameterization in equation 14, the gradient with respect to the network parameters $\bm{\theta}$ becomes
$\frac{dL}{d\bm{\theta}}=\frac{dL}{d\mathcal{P}}\frac{d\mathcal{P}}{d\mathcal{N}_{\bm{\theta}}}\frac{d\mathcal{N}_{\bm{\theta}}}{d\bm{\theta}}=\tilde{\mathcal{P}}\frac{dL}{d\mathcal{P}}\frac{d\mathcal{N}_{\bm{\theta}}}{d\bm{\theta}}.$ (17)
Thus, the core physical gradient $\frac{dL}{d\mathcal{P}}$ is consistently inherited across both stages, ensuring smooth convergence and stable optimization dynamics.
Neural network architecture and problem-specific design The feed-forward neural networks in the second stage are adaptively designed, with the number of residual blocks and neurons per layer customized for each application. Specifically, in the thermal battery problem, the scalar references for thermal conductivity in different directions ($\tilde{k}_{x}$, $\tilde{k}_{z}$) and heat capacity ($\tilde{C_{p}}$) are first determined. In the cell migration problem, three scalar parameters, $\tilde{\gamma}$, $\tilde{\lambda_{1}}$, and $\tilde{\lambda_{2}}$, are similarly estimated. In both cases, the neural network models state-dependent field parameters based on a single state variable, with a network input size of 1 (e.g., temperature $T$ or density $\rho$, rescaled approximately from -1 to 1). A compact network with 20 neurons per hidden layer and two residual blocks is used to ensure efficient training.
In contrast, for the flow and cardiac applications, the conductivity parameters $K$ or $D$ depend on a 2D space, requiring larger neural networks: 50 neurons with 3 residual blocks for the flow problem, and 150 neurons with 4 blocks for the cardiac problem. The spatial inputs ($x$, $y$) are rescaled from 1-1 to $1$ in both dimensions. Notably, in the flow problem, multiple parameter distributions can satisfy the same setup due to the nature of equation 7, which involves taking spatial derivatives. Therefore, the first-stage scalar estimation is omitted, and the neural network is directly trained to predict relative variations by outputting $\mathcal{N}_{\bm{\theta}}+1\times 10^{-3}$. An absolute value activation with a small offset is applied at the final layer to ensure strictly positive predictions and avoid physically meaningless zeros.
Additionally, Supplementary Note 4 presents an ablation study evaluating the impact of neural network size on parameter estimation performance, while Supplementary Note 5 provides a detailed comparison of this training strategy against alternatives, including direct field parameter learning and using a single neural network for all parameters.
Loss function and model training Neptune relies solely on observations from physical phenomena for training. The loss function is defined purely using the mean absolute error (MAE):
$L=\frac{1}{N}\sum_{i=1}^{N}|\mathcal{S}_{i}-\hat{\mathcal{S}}_{i}|$ (18)
where $\mathcal{S}$ and $\hat{\mathcal{S}}$ represent the ground truth (observations) and predicted responses, respectively, $N$ is the number of points. Preliminary experiments showed that mean squared error (MSE) yielded similar performance, and thus, MAE was chosen for simplicity. Regularization terms (e.g., weight decay) were also tested but adversely affected optimization and were therefore omitted.
During training, observations are split into mini-batches of size 16 (reduced to 8 for the cell migration problem due to limited samples). In the first stage, the loss is minimized to directly learn $\tilde{\mathcal{P}}$, with the learning rate adaptively decreasing from 1 to $10^{-2}$ until convergence, where $\tilde{\mathcal{P}}$ stabilizes without further change. Dashed lines in the figure illustrate the gradient flow during optimization. In the second stage, the loss is minimized to optimize the neural network weights $\bm{\theta}$, with a smaller learning rate ranging from $10^{-4}$ to $10^{-2}$ depending on network size. The best-performing $\bm{\theta}$, corresponding to the lowest loss value, is saved as the final trained model. The field parameter is then predicted using equation 14 with the corresponding spatial or state inputs.
To update the learnable weights ($\tilde{\mathcal{P}}$ or $\bm{\theta}$), Neptune employs gradient-based optimization through adjoint sensitivity analysis [62], leveraging advancements in neural ODEs and reverse-mode automatic differentiation [68, 63]. For the scalar parameters $\tilde{\mathbf{p}}$, the adjoint method directly computes the loss gradient through the PDE solver, as shown in equation 16. For the neural network weights $\bm{\theta}$, an additional application of the chain rule through the neural network is required, as detailed in equation 17.
Importantly, Neptune differs fundamentally from traditional neural ODE approaches: instead of modeling unknown or partially known governing expressions with neural networks, it tackles complex PDEs with fully known equations, while neural networks serve only as correction terms for modeling unknown field parameters. To clarify the underlying optimization, we next emphasize the derivation of the common term $\frac{dL}{d\mathcal{P}}$ for a standard differential equation. The adjoint method requires defining a scalar functional $g(u,p)$, where $u(t,p)$ is the numerical solution to the differential equation $\frac{d\mathcal{S}(t,\mathcal{P})}{dt}=f(t,\mathcal{S},\mathcal{P})$ with $t\in[0,T]$ and initial condition $\mathcal{S}(t_{0},\mathcal{P})=\mathcal{S}_{0}$. In our experiments, we assume a known initial condition.
In our work, though the PDE system is discretized spatially using equation 13, this semi-discretization does not fundamentally alter the adjoint sensitivity formulation. Adjoint sensitivity analysis is still applied to compute the gradient of $g$ with respect to the parameters $\mathcal{P}$:
$L(\mathcal{S},\mathcal{P})=\int_{t_{0}}^{T}g(\mathcal{S}(t,\mathcal{P}),\mathcal{P})\,dt$ (19)
To derive the adjoint equation [105], we introduce the Lagrange multiplier $\bm{\lambda}$ to form:
$I(\mathcal{P})=L(\mathcal{P})-\int_{t_{0}}^{T}\bm{\lambda}^{\top}\left(\frac{d\mathcal{S}}{dt}-f(t,\mathcal{S},\mathcal{P})\right)\,dt$ (20)
Since $\mathcal{S}^{\prime}=f(\mathcal{S},\mathcal{P},t)$, we have that:
$\frac{dL}{d\mathcal{P}}=\frac{dI}{d\mathcal{P}}$ $=\int_{t_{0}}^{T}\left(\frac{\partial g}{\partial\mathcal{P}}+\frac{\partial g}{\partial\mathcal{S}}\frac{\partial\mathcal{S}}{\partial\mathcal{P}}\right)dt$
$\quad-\int_{t_{0}}^{T}\bm{\lambda}^{\top}\left(\frac{d}{dt}\left(\frac{\partial\mathcal{S}}{\partial\mathcal{P}}\right)-\frac{\partial f}{\partial\mathcal{S}}\frac{\partial\mathcal{S}}{\partial\mathcal{P}}-\frac{\partial f}{\partial\mathcal{P}}\right)dt,$ (21)
where $\frac{\partial\mathcal{S}}{\partial\mathcal{P}}$ represents the sensitivity of the state variables to the parameters. After applying integration by parts to the term involving $\bm{\lambda}^{\top}\frac{d}{dt}\left(\frac{\partial\mathcal{S}}{\partial\mathcal{P}}\right)$, we require that:
$\frac{d\bm{\lambda}(t)}{dt}=\frac{\partial g(\mathcal{S}(t,\mathcal{P}),\mathcal{P})}{\partial\mathcal{S}}$ $-\bm{\lambda}(t)\frac{\partial f(t,\mathcal{S}(t,\mathcal{P}),\mathcal{P})}{\partial\mathcal{S}},$
$\bm{\lambda}(T)$ $=0$ (22)
where $\frac{\partial f}{\partial\mathcal{S}}$ is the Jacobian of the system with respect to the state $\mathcal{S}$, and $\frac{\partial f}{\partial\mathcal{P}}$ is the Jacobian with respect to the parameters. The solution to the adjoint problem provides the sensitivities through the integral:
$\frac{dL}{d\mathcal{P}}=\int_{t_{0}}^{T}\left(\bm{\lambda}^{\top}\frac{\partial f}{\partial\mathcal{P}}+\frac{\partial g}{\partial\mathcal{P}}\right)dt+\bm{\lambda}^{\top}(t_{0})\frac{\partial\mathcal{S}(t_{0})}{\partial\mathcal{P}}$ (23)
To compute the term $\frac{\partial L}{\partial\mathcal{P}}$, the adjoint problem is formulated and solved as an ODE system. This is achieved using a specialized sensitivity analysis framework [68], which provides a range of numerical methods tailored to different types of PDE systems.
Data generation For the cell migration and proliferation problem, experimental data are sourced from previous research [99]. For other data, we generate data by employing FDM to solve the PDEs in each problem. For battery thermal runaway modeling, the mesh size is set to $8$ in each spatial dimension, with a timestep of $200\,\text{s}$ (i.e., $100$ minutes over $30$ steps). The two state variables, temperature $T$ and concentration $c$, are updated simultaneously at each step, along with the temperature-dependent parameters. Secondly, in the flow problem, the PDEs are solved in multiple stages. The mesh size is set to $22\times 22$, and a reference parameter $K$, based on previous research [17], is regenerated. The hydraulic head $h$ is computed using Darcy’s law (equation 7), followed by the velocity fields in the $x$ and $y$ directions, solved using $K$ and $h$ (equation 8). The resulting velocity field is then used to solve the time-dependent concentration equation for $u$ (equation 6). Thirdly, in the cardiac electrophysiology problem, the mesh size is set to $50$ in both $x$ and $y$ dimensions. The two state variables, $V$ and $W$, are computed at each timestep (1TUTU) during the simulation. Lastly, in the cell migration and proliferation problem, the mesh size is fixed at $38$ points, corresponding to the spatial measurement positions. Due to varying gaps between the first two and the last two points, the standard FDM is modified to accommodate a non-uniform finite difference mesh. There are $5$ timesteps: the first for setting initial conditions, and the remaining $4$ for training.
Notably, Supplementary Note 3 expands on the mathematical derivation of finite difference discretization [106, 107, 108], explains the treatment of non-uniform meshes at geometry boundaries for the cell migration and proliferation problem, and discusses limitations of FDM. The training data, consisting of only the state responses, is generated numerically using the PDE solver for the first three scenarios, with different levels of Gaussian noise added to ensure robustness (details in Supplementary Note 1).
Author contributions N.L. and V.B. supervised the study. X.L., N.L., and V.B. conceived the initial idea. X.L. and M.M. conducted the numerical experiments. X.L., M.M., N.L., and V.B. wrote the paper. X.L., M.M., and R.G. contributed to the supplementary.
Competing interests The authors declare no competing interests.
Correspondence and requests for materials should be addressed to vishnu@msu.edu and lajnefni@msu.edu.